# Deviation Inequalities for Rényi Divergence Estimators via Variational Expression

### v1

@@ 4.1. Proof of Theorem 1
$\check{\zeta}_{n,d,\tau}(z):=n^{\frac{d+12-8\tau}{d+12}}z^{2}\wedge\frac{n^{\frac{d+12-4\tau}{d+12}}z}{\log(n+1)}.$
@@
where
$\hat{\zeta}_{n,d,p,\tau}(z):=n^{\frac{d+12-8\tau}{d+12}}z^{2}\wedge\frac{n^{\frac{d+12-4\tau}{d+12}}z}{\log(n+1)}\wedge\frac{n^{\frac{p\tau}{d+12}}}{\log(n+1)}.$ (43)
Taking supremum over all $\mu,\nu\in\mathsf{SG}(L)$ such that $\left\|X\right\|_{\psi_{p}}\leq L$ yields (9), thus completing the proof.
@@ 4.2. Proof of Theorem 2
$\qquad\qquad\qquad\qquad\leq e^{-\left(\frac{nz^{2}}{8\left\|F_{n}*\varphi_{\sigma}\right\|_{2,\mu}^{2}}\wedge\frac{nz}{22\left\|F_{n}*\varphi_{\sigma}\right\|_{\infty,\mu}}\right)}+e^{-\left(\frac{nz^{2}}{8\left\|e^{F_{n}}*\varphi_{\sigma}\right\|_{2,\mu}^{2}}\wedge\frac{nz}{22\left\|e^{F_{n}}*\varphi_{\sigma}\right\|_{\infty,\nu}}\right)}.$ (45)
The proof of Lemma 5 follows similar to Lemma 2 up to (66), and then applies (12) along with union bound to obtain (44), and (85) with $\eta=1+\sqrt{6}$ to derive (45).
Theorem 2 follows from (45) by noting that for $\mu,\nu\in\mathcal{P}\left(\mathbb{B}_{d}(r)\right)$, and function classes $\mathcal{F}_{\sigma}(b_{r,\sigma})$ and $\mathcal{E}_{\sigma}(b_{r,\sigma})$ with envelopes $F_{b_{r,\sigma}}*\varphi_{\sigma}$ and $e^{F_{b_{r,\sigma}}}*\varphi_{\sigma}$, respectively, we have
@@ 4.3. Proof of Theorem 3
where the penultimate inequality used $\left|\log\frac{x}{y}\right|\leq\frac{|x-y|}{x\wedge y}$.
Note that since $\mathrm{spt}(\mu)\subseteq\mathbb{B}_{d}(r)$, we have
@@
where we used that $\mathbb{E}[\left\|Z\right\|]=\sqrt{d}\sigma$ for $Z\sim N(0,\sigma^{2}I_{d})$ and $\kappa_{\alpha,r,d,\sigma}:=e^{\frac{\alpha b_{r,\sigma}(1+r+\sqrt{d}\sigma)}{2}}$. Similarly, using $\mathrm{spt}(\hat{\mu}_{n}),\mathrm{spt}(\hat{\nu}_{n}),\mathrm{spt}(\nu)\subseteq\mathbb{B}_{d}(r)$, we can lower bound the terms in the denominator of the expression in (47), which upon substitution therein yields
@@
$=\sup_{f\in\bar{\mathcal{E}}_{\sigma,\alpha,r,d}}\left|(\hat{\mu}_{n}-\mu)(f)\right|+\sup_{f\in\tilde{\mathcal{E}}_{\sigma,\alpha,r,d}}\left|(\hat{\nu}_{n}-\nu)(f)\right|.$
@@
$\bar{\mathcal{E}}_{\sigma,\alpha,r,d}$ $:=\left\{\frac{\alpha\kappa_{|\alpha-1|,r,d,\sigma}}{\left|\alpha-1\right|}e^{|\alpha-1|f}*\varphi_{\sigma}:f\in\mathcal{F}(b_{r,\sigma})\right\},$
$\tilde{\mathcal{E}}_{\sigma,\alpha,r,d}$ $:=\left\{\kappa_{\alpha,r,d,\sigma}e^{\alpha f}*\varphi_{\sigma}:f\in\mathcal{F}(b_{r,\sigma})\right\},$
@@
$\bar{E}_{\sigma,\alpha,r,d}$ $:=\frac{\alpha\kappa_{|\alpha-1|,r,d,\sigma}}{\left|\alpha-1\right|}e^{|\alpha-1|F_{b_{r,\sigma}}},$
$\tilde{E}_{\sigma,\alpha,r,d}$ $:=\kappa_{\alpha,r,d,\sigma}e^{\alpha F_{b_{r,\sigma}}}.$
@@
$J(1,\bar{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma},\bar{\mathcal{E}}_{\sigma,\alpha,r,d},\mu)\vee J(1,\tilde{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma},\tilde{\mathcal{E}}_{\sigma,\alpha,r,d},\nu)$ $\leq(3ed)^{3(d+2)}(r+1)^{\frac{d}{2}+1}\sigma^{-\frac{d}{2}}=:\tilde{\xi}_{r,d,\sigma},$
@@
$t_{n,\sigma,\mu,\nu}$
$:=J(1,\bar{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma},\bar{\mathcal{E}}_{\sigma,\alpha,r,d},\mu)\left\|\bar{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma}\right\|_{2,\mu}+J(1,\tilde{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma},\tilde{\mathcal{E}}_{\sigma,\alpha,r,d},\nu)\left\|\tilde{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma}\right\|_{2,\nu}$
$\lesssim(3ed)^{3(d+2)}(r+1)^{\frac{d}{2}+1}\sigma^{-\frac{d}{2}}\left(\frac{\alpha}{\left|\alpha-1\right|}e^{|\alpha-1|b_{r,\sigma}\left(1+r+\frac{\sqrt{d}\sigma}{2}+d|\alpha-1|b_{r,\sigma}\sigma^{2}\right)}+e^{\alpha b_{r,\sigma}\left(1+r+\frac{\sqrt{d}\sigma}{2}+d\alpha b_{r,\sigma}\sigma^{2}\right)}\right)$
$\lesssim\tilde{\xi}_{r,d,\sigma}\lambda_{\alpha,d,r,\sigma},$
@@ 4.4. Proof of Corollary 1
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq e^{-\left(nz^{2}\wedge nz\right)},$
@@ 4.5. Proof of Proposition 1
The inequality (15a) follows via same steps as in [67, Lemma 3] up to replacing Lebesgue measure by $\rho$ and taking $p_{\mu}$, $p_{\nu}$ as densities w.r.t. $\rho$. We next show (15b) using the same techniques therein. For $\tau\in[0,1]$, let $z_{\mu,\tau}(x):=(1-\tau)p_{\mu}(x)+\tau\mu*\varphi_{\sigma}(x)$ and $z_{\nu,\tau}(x):=(1-\tau)p_{\nu}(x)+\tau\nu*\varphi_{\sigma}(x)$. By Taylor’s theorem,
@@
$\quad+(1-\alpha)\big{(}\nu*\varphi_{\sigma}(x)-p_{\nu}(x)\big{)}\int_{0}^{1}\bigg{(}\frac{z_{\mu,\tau}(x)}{z_{\nu,\tau}(x)}\bigg{)}^{\alpha}d\tau.$ (49)
@@
which follows under the Lipschitz assumption on the densities similarly to the proof of [67, Lemma 3]. On the other hand, for $\alpha\in(1,\infty)$, integrating w.r.t. $\rho$ in (49) leads to
@@
$e^{(\alpha-1)\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)}=\int p_{\mu}(x)^{\alpha}p_{\nu}(x)^{1-\alpha}d\rho\geq M^{-\alpha},$ (53a)
and similarly
$e^{(\alpha-1)\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)}\geq M^{-\alpha}.$ (53b)
@@ 4.6. Proof of Theorem 4
Suppose $\mathcal{F}$ satisfies Assumption 1 with envelope $F\in L^{p}(\mu)$ for some $p\geq 2$. For $X^{n}\sim\mu^{\otimes n}$, let $Z_{n}$ be as given in (11a) (with $\mathcal{F}_{n}=\mathcal{F}$ for all $n\in\mathbb{N}$). Then
@@
$\bar{\mathcal{G}}_{b,\alpha,\beta,\rho}$ $:=\left\{\frac{\alpha e^{\left|\alpha-1\right|b}}{\left|\alpha-1\right|}e^{(\alpha-1)g}:g\in\mathcal{G}_{b,\beta,\rho}\right\},$ (58a)
$\tilde{\mathcal{G}}_{b,\alpha,\beta,\rho}$ $:=\big{\{}e^{\alpha(b+g)}:g\in\mathcal{G}_{b,\beta,\rho}\big{\}}.$ (58b)
@@
$N\Big{(}\alpha e^{2\left|\alpha-1\right|b}\epsilon,\bar{\mathcal{G}}_{b,\alpha,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)}\leq N\Big{(}\epsilon,\mathcal{G}_{b,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)},$ (59a)
$N\Big{(}\alpha e^{2\alpha b}\epsilon,\tilde{\mathcal{G}}_{b,\alpha,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)}\leq N\Big{(}\epsilon,\mathcal{G}_{b,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)}.$ (59b)
@@
$\sqrt{V_{n}}$ $\leq\sqrt{n}\frac{\alpha e^{2\left|\alpha-1\right|b}}{\left|\alpha-1\right|}+\sqrt{2\frac{\alpha e^{2\left|\alpha-1\right|b}}{\left|\alpha-1\right|}\mathbb{E}\big{[}Z_{n,\bar{\mathcal{G}}_{b,\alpha,\beta,\rho}}\big{]}}$
$\leq\sqrt{n}\frac{\alpha e^{2\left|\alpha-1\right|b}}{\left|\alpha-1\right|}+\frac{\sqrt{2c_{d,\beta}}\alpha e^{2\left|\alpha-1\right|b}b^{\frac{a}{2}}n^{\frac{1}{4}}}{\left|\alpha-1\right|^{1-\frac{a}{2}}}$
@@
Applying (62b), it follows similar to (57) that for $z\geq 0$:
@@ 5.1. Proof of Lemma 2
To handle the case of compactly supported distributions, we will use a version of Talagrand’s concentration inequality given in [12, Theorem 2.3].
@@
Suppose that $\mathcal{F}$ satisfies Assumption 1 and $\left\|f\right\|_{\infty,\mu}\leq b$ for all $f\in\mathcal{F}$. For $X^{n}\sim\mu^{\otimes n}$, let $Z_{n}$ be as given in (11a) (with $\mathcal{F}_{n}=\mathcal{F}$ for all $n\in\mathbb{N}$). Let $\theta^{2}:=\sup_{f\in\mathcal{F}}\mathbb{E}_{\mu}\big{[}\big{(}f-\mathbb{E}_{\mu}[f]\big{)}^{2}\big{]}$ and $V_{n}:=n\theta^{2}+2b\mathbb{E}[Z_{n}]$. Then, for all $z\geq 0$,
@@
To show (24), we will use the following Fuk-Nagaev type inequality which is applicable for a sequence of function classes $(\mathcal{F}_{n})_{n\in\mathbb{N}}$ with unbounded envelope. It is a restatement of [2, Theorem 4] in a form with explicit constants which is useful for our purposes.
@@
In the setting of Lemma 1, for $c_{q}$ as specified in (83),
$\mathbb{P}\Big{(}Z_{n}\geq 2\mathbb{E}[Z_{n}]+z\Big{)}\leq e^{-\big{(}\frac{z^{2}}{128n\left\|F_{n}\right\|_{2,\mu}^{2}}\wedge\frac{3z}{256\bar{M}_{n}}\big{)}}+3e^{-\left(\frac{z}{c_{q}\left\|M_{n}\right\|_{\psi_{q}}}\right)^{q}}.$ (63)
The constants in Theorem 7 are determined similar to that in Lemma 1.
Proceeding with the proof of (24), let $X^{n}\sim\mu^{\otimes n}$, $Y^{n}\sim\nu^{\otimes n}$, and $\mathcal{F}_{n}$ be the function class having the properties specified in Lemma 2. Recall that $f_{\mu*\gamma_{\sigma},\nu*\gamma_{\sigma}}^{\star}:=\log\left(\mu*\varphi_{\sigma}/\nu*\varphi_{\sigma}\right)$. Since $f_{\hat{\mu}_{n}*\gamma_{\sigma},\hat{\nu}_{n}*\gamma_{\sigma}}^{\star},f_{\mu*\gamma_{\sigma},\nu*\gamma_{\sigma}}^{\star}\in\mathcal{F}_{n}$ by assumption,
$\sup_{f\in\mathcal{F}_{n}}\mathbb{E}_{\hat{\mu}_{n}*\gamma_{\sigma}}[f]-\mathbb{E}_{\hat{\nu}_{n}*\gamma_{\sigma}}[e^{f}]+1$ $\leq\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)$
$=\mathbb{E}_{\hat{\mu}_{n}*\gamma_{\sigma}}\big{[}f_{\hat{\mu}_{n}*\gamma_{\sigma},\hat{\nu}_{n}*\gamma_{\sigma}}^{\star}\big{]}-\mathbb{E}_{\hat{\nu}_{n}*\gamma_{\sigma}}\big{[}e^{f_{\hat{\mu}_{n}*\gamma_{\sigma},\hat{\nu}_{n}*\gamma_{\sigma}}^{\star}}\big{]}+1,$
$\sup_{f\in\mathcal{F}_{n}}\mathbb{E}_{\mu*\gamma_{\sigma}}[f]-\mathbb{E}_{\nu*\gamma_{\sigma}}[e^{f}]+1$ $\leq\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)$
$=\mathbb{E}_{\mu*\gamma_{\sigma}}\big{[}f_{\mu*\gamma_{\sigma},\nu*\gamma_{\sigma}}^{\star}\big{]}-\mathbb{E}_{\nu*\gamma_{\sigma}}\big{[}e^{f_{\mu*\gamma_{\sigma},\nu*\gamma_{\sigma}}^{\star}}\big{]}+1.$
@@
$\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)=\sup_{f\in\mathcal{F}_{n}}\mathbb{E}_{\hat{\mu}_{n}*\gamma_{\sigma}}[f]-\mathbb{E}_{\hat{\nu}_{n}*\gamma_{\sigma}}[e^{f}]+1,$ (64a)
$\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)=\sup_{f\in\mathcal{F}_{n}}\mathbb{E}_{\mu*\gamma_{\sigma}}[f]-\mathbb{E}_{\nu*\gamma_{\sigma}}[e^{f}]+1.$ (64b)
Moreover, since $\left|f_{\mu*\gamma_{\sigma},\nu*\gamma_{\sigma}}^{\star}(x)\right|\leq F_{n}(x)$ for all $x\in\mathbb{R}^{d}$ and $\left\|F_{n}*\varphi_{\sigma}(X)\right\|_{\psi_{q}}<\infty$ for some $q\in(0,1]$ implies that $\left\|F_{n}\right\|_{1,\mu*\gamma_{\sigma}}=\left\|F_{n}*\varphi_{\sigma}\right\|_{1,\mu}<\infty$ by (103) given below, $\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)<\infty$. Then, from (46) and (64), we obtain
$\big{|}\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\gamma_{\sigma}\middle\|\hat{\nu}_{n}*\gamma_{\sigma}\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)\big{|}$
$\leq\sup_{f\in\mathcal{F}_{n}}\left|(\hat{\mu}_{n}*\varphi_{\sigma}-\mu*\varphi_{\sigma})(f)-(\hat{\nu}_{n}*\varphi_{\sigma}-\nu*\varphi_{\sigma})(e^{f})\right|$
$\leq\sup_{f\in\mathcal{F}_{n}}\left|(\hat{\mu}_{n}*\varphi_{\sigma}-\mu*\varphi_{\sigma})(f)\right|+\sup_{f\in\mathcal{F}_{n}}\left|(\hat{\nu}_{n}*\varphi_{\sigma}-\nu*\varphi_{\sigma})(e^{f})\right|$
$=\frac{Z_{n,\mathcal{F}_{n,\sigma}}+\tilde{Z}_{n,\mathcal{E}_{n,\sigma}}}{n},$

### v3

@@ 4.1. Proof of Theorem 1
$\check{\zeta}_{n,d,\tau}(z):=n^{\frac{d+12-8\tau}{d+12}}z^{2}\wedge n^{\frac{d+12-4\tau}{d+12}}z.$
@@
where $\hat{\zeta}_{n,d,p,\tau}(z)$ is as given in (10). Taking supremum over all $\mu,\nu\in\mathsf{SG}(L)$ such that $\left\|X\right\|_{\psi_{p}}\leq L$ yields (9), thus completing the proof.
@@ 4.2. Proof of Theorem 2
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+2(16)^{p}\left(\left\|F_{n}*\varphi_{\sigma}\right\|_{p,\mu}^{p}+\left\|e^{F_{n}}*\varphi_{\sigma}\right\|_{p,\nu}^{p}\right)n^{1-p}.$ (42)
@@
$\qquad\qquad\qquad\leq e^{-\left(\frac{nz^{2}}{8\left\|F_{n}*\varphi_{\sigma}\right\|_{2,\mu}^{2}}\wedge\frac{nz}{22\left\|F_{n}*\varphi_{\sigma}\right\|_{\infty,\mu}}\right)}+e^{-\left(\frac{nz^{2}}{8\left\|e^{F_{n}}*\varphi_{\sigma}\right\|_{2,\mu}^{2}}\wedge\frac{nz}{22\left\|e^{F_{n}}*\varphi_{\sigma}\right\|_{\infty,\nu}}\right)}.$ (43)
The proof of Lemma 5 follows similar to Lemma 2 up to (61), and then applies (14) along with $\overline{M_{n}}\leq\left\|M_{n}(X^{n})\right\|_{p,\mu}\leq n^{\frac{1}{p}}\left\|F_{n}\right\|_{p,\mu}$ to obtain (42), and (79) with $\eta=1+\sqrt{6}$ to derive (43).
Theorem 2 follows from (43) by noting that for $\mu,\nu\in\mathcal{P}\left(\mathbb{B}_{d}(r)\right)$, and function classes $\mathcal{F}_{\varphi_{\sigma}}(b_{r,\sigma})$ and $\mathcal{E}_{\varphi_{\sigma}}(b_{r,\sigma})$ with envelopes $F_{b_{r,\sigma}}*\varphi_{\sigma}$ and $e^{F_{b_{r,\sigma}}}*\varphi_{\sigma}$, respectively, we have
@@ 4.3. Proof of Theorem 3
where the penultimate inequality used $\left|\log\frac{x}{y}\right|\leq\frac{|x-y|}{x\wedge y}$. Note that since $\mathrm{spt}(\mu)\subseteq\mathbb{B}_{d}(r)$, we have
@@
[…]
@@
$=\sup_{f\in\bar{\mathcal{E}}_{\varphi_{\sigma},\alpha,r,d}}\left|(\hat{\mu}_{n}-\mu)(f)\right|+\sup_{f\in\tilde{\mathcal{E}}_{\varphi_{\sigma},\alpha,r,d}}\left|(\hat{\nu}_{n}-\nu)(f)\right|.$
@@
$\bar{\mathcal{E}}_{\varphi_{\sigma},\alpha,r,d}$ $:=\left\{\frac{\alpha\kappa_{|\alpha-1|,r,d,\sigma}}{\left|\alpha-1\right|}e^{|\alpha-1|f}*\varphi_{\sigma}:f\in\mathcal{F}(b_{r,\sigma})\right\}$
$\mbox{and}\quad\tilde{\mathcal{E}}_{\varphi_{\sigma},\alpha,r,d}$ $:=\left\{\kappa_{\alpha,r,d,\sigma}e^{\alpha f}*\varphi_{\sigma}:f\in\mathcal{F}(b_{r,\sigma})\right\},$
@@
$\bar{E}_{\sigma,\alpha,r,d}$ $:=\frac{\alpha\kappa_{|\alpha-1|,r,d,\sigma}}{\left|\alpha-1\right|}e^{|\alpha-1|F_{b_{r,\sigma}}}\quad\mbox{and}\quad\tilde{E}_{\sigma,\alpha,r,d}:=\kappa_{\alpha,r,d,\sigma}e^{\alpha F_{b_{r,\sigma}}}.$
@@
$J(1,\bar{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma},\bar{\mathcal{E}}_{\varphi_{\sigma},\alpha,r,d},\mu)\vee J(1,\tilde{E}_{\sigma,\alpha,r,d}*\varphi_{\sigma},\tilde{\mathcal{E}}_{\varphi_{\sigma},\alpha,r,d},\nu)$ $\lesssim\tilde{\xi}_{r,d,\sigma},$
@@
$t_{n,\varphi_{\sigma},\mu,\nu}$
[…]
$\lesssim\tilde{\xi}_{r,d,\sigma}\left(\frac{\alpha}{\left|\alpha-1\right|}e^{|\alpha-1|b_{r,\sigma}\left(1+r+\frac{\sqrt{d}\sigma}{2}+d|\alpha-1|b_{r,\sigma}\sigma^{2}\right)}+e^{\alpha b_{r,\sigma}\left(1+r+\frac{\sqrt{d}\sigma}{2}+d\alpha b_{r,\sigma}\sigma^{2}\right)}\right)$
[…]
@@ 4.4. Proof of Corollary 1
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\leq e^{-\left(nz^{2}\wedge nz\right)},$
@@ 4.5. Proof of Proposition 1
The inequality (17a) follows via same steps as in [76, Lemma 3] up to replacing Lebesgue measure by $\rho$ and taking $p_{\mu}$, $p_{\nu}$ as densities w.r.t. $\rho$. We next show (17b) using the same techniques therein. For $\tau\in[0,1]$, let $z_{\mu,\tau}(x):=(1-\tau)p_{\mu}(x)+\tau\mu*\varphi_{\sigma}(x)$ and $z_{\nu,\tau}(x):=(1-\tau)p_{\nu}(x)+\tau\nu*\varphi_{\sigma}(x)$. By Taylor’s theorem,
@@
$\qquad\qquad+(1-\alpha)\big{(}\nu*\varphi_{\sigma}(x)-p_{\nu}(x)\big{)}\int_{0}^{1}\bigg{(}\frac{z_{\mu,\tau}(x)}{z_{\nu,\tau}(x)}\bigg{)}^{\alpha}d\tau.$ (47)
@@
which follows under the Lipschitz assumption on the densities similarly to the proof of [76, Lemma 3]. On the other hand, for $\alpha\in(1,\infty)$, integrating w.r.t. $\rho$ in (47) leads to
@@
$e^{(\alpha-1)\mathsf{D}_{\alpha}\left(\mu\middle\|\nu\right)}=\int p_{\mu}(x)^{\alpha}p_{\nu}(x)^{1-\alpha}d\rho\geq M^{-\alpha}\quad\mbox{ and }\quad e^{(\alpha-1)\mathsf{D}_{\alpha}\left(\mu*\gamma_{\sigma}\middle\|\nu*\gamma_{\sigma}\right)}\geq M^{-\alpha}.$ (51)
@@ 4.6. Proof of Theorem 4
Suppose $\mathcal{F}$ satisfies Assumption 1 with envelope $F\in L^{p}(\mu)$ for some $p\geq 2$. For $X^{n}\sim\mu^{\otimes n}$, let $Z_{n}$ be as given in (12) (with $\mathcal{F}_{n}=\mathcal{F}$ for all $n\in\mathbb{N}$). Then
@@
$\bar{\mathcal{G}}_{b,\alpha,\beta,\rho}$ $:=\left\{\frac{\alpha e^{\left|\alpha-1\right|b}}{\left|\alpha-1\right|}e^{(\alpha-1)g}:g\in\mathcal{G}_{b,\beta,\rho}\right\}\quad\mbox{and}\quad\tilde{\mathcal{G}}_{b,\alpha,\beta,\rho}:=\big{\{}e^{\alpha(b+g)}:g\in\mathcal{G}_{b,\beta,\rho}\big{\}}.$
@@
$N\Big{(}\alpha e^{2\left|\alpha-1\right|b}\epsilon,\bar{\mathcal{G}}_{b,\alpha,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)}$ $\leq N\Big{(}\epsilon,\mathcal{G}_{b,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)},$
$\mbox{and}\quad N\Big{(}\alpha e^{2\alpha b}\epsilon,\tilde{\mathcal{G}}_{b,\alpha,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)}$ $\leq N\Big{(}\epsilon,\mathcal{G}_{b,\beta,\rho},\left\|\cdot\right\|_{2,\gamma}\Big{)}.$
@@
$\sqrt{V_{n}}\leq\sqrt{n}\frac{\alpha e^{2\left|\alpha-1\right|b}}{\left|\alpha-1\right|}+\sqrt{2\frac{\alpha e^{2\left|\alpha-1\right|b}}{\left|\alpha-1\right|}\mathbb{E}\big{[}Z_{n,\bar{\mathcal{G}}_{b,\alpha,\beta,\rho}}\big{]}}$ $\leq\sqrt{n}\frac{\alpha e^{2\left|\alpha-1\right|b}}{\left|\alpha-1\right|}+\frac{\sqrt{2c_{d,\beta}}\alpha e^{2\left|\alpha-1\right|b}b^{\frac{a}{2}}n^{\frac{1}{4}}}{\left|\alpha-1\right|^{1-\frac{a}{2}}}$
@@
Applying (58b), it follows similar to (55) that for $z\geq 0\mspace${2.0mu}:
@@ 5.1. Proof of Lemma 2
To handle the case of compactly supported distributions, we will use a version of Talagrand’s concentration inequality given in [15, Theorem 2.3].
@@
Suppose that $\mathcal{F}$ satisfies Assumption 1 and $\left\|f\right\|_{\infty,\mu}\leq b$ for all $f\in\mathcal{F}$. For $X^{n}\sim\mu^{\otimes n}$, let $Z_{n}$ be as given in (12) (with $\mathcal{F}_{n}=\mathcal{F}$ for all $n\in\mathbb{N}$). Let $\theta^{2}:=\sup_{f\in\mathcal{F}}\mathbb{E}_{\mu}\big{[}\big{(}f-\mathbb{E}_{\mu}[f]\big{)}^{2}\big{]}$ and $V_{n}:=n\theta^{2}+2b\mathbb{E}[Z_{n}]$. Then, for all $z\geq 0$,
@@
To show (25), we will use the following Fuk-Nagaev type inequality which is applicable for a sequence of function classes $(\mathcal{F}_{n})_{n\in\mathbb{N}}$ with unbounded envelope. It is a restatement of [3, Theorem 2] in a form with explicit constants which is useful for our purposes.
@@
In the setting of Lemma 1, if $F_{n}$ satisfies the weaker integrability condition $\left\|F_{n}\right\|_{p,\mu}<\infty$ for some $p\geq 2$ (instead of $\left\|F_{n}\right\|_{\psi_{q}}<\infty$ for some $q\in(0,1]$),
$\mathbb{P}\Big{(}Z_{n}\geq 2\mathbb{E}[Z_{n}]+z\Big{)}\leq e^{-\big{(}\frac{z^{2}}{128n\left\|F_{n}\right\|_{2,\mu}^{2}}\wedge\frac{3z}{256\overline{M_{n}}}\big{)}}+2(16)^{p}\mathbb{E}\big{[}\big{(}M_{n}(X^{n})\big{)}^{p}\big{]}z^{-p},$ (59)
where $\overline{M_{n}}\leq\left\|M_{n}(X^{n})\right\|_{p,\mu}\leq n^{\frac{1}{p}}\left\|F_{n}\right\|_{p,\mu}$.
The constants in Theorem 7 are determined similar to that in Lemma 1, and the upper bound on $\left\|M_{n}(X^{n})\right\|_{p,\mu}$ follows due to the maximal inequality for Orlicz norm with $\psi(x)=x^{p}$ (see equation above [83, Lemma 2.2.2]).
Proceeding with the proof of (25), let $X^{n}\sim\mu^{\otimes n}$, $Y^{n}\sim\nu^{\otimes n}$, and $\mathcal{F}_{n}$ be the function class having the properties specified in Lemma 2. Recall that $f_{\mu*\Phi,\nu*\Phi}^{\star}:=\log\left(\mu*\phi/\nu*\phi\right)$. Since $f_{\hat{\mu}_{n}*\Phi,\hat{\nu}_{n}*\Phi}^{\star},f_{\mu*\Phi,\nu*\Phi}^{\star}\in\mathcal{F}_{n}$ by assumption,
[…]
$\sup_{f\in\mathcal{F}_{n}}\mathbb{E}_{\mu*\Phi}[f]-\mathbb{E}_{\nu*\Phi}[e^{f}]+1$ $\leq\mathsf{D}_{\mathsf{KL}}\left(\mu*\Phi\middle\|\nu*\Phi\right)=\mathbb{E}_{\mu*\Phi}\big{[}f_{\mu*\Phi,\nu*\Phi}^{\star}\big{]}-\mathbb{E}_{\nu*\Phi}\big{[}e^{f_{\mu*\Phi,\nu*\Phi}^{\star}}\big{]}+1.$
@@
$\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\Phi\middle\|\hat{\nu}_{n}*\Phi\right)=\sup_{f\in\mathcal{F}_{n}}\mathbb{E}_{\hat{\mu}_{n}*\Phi}[f]-\mathbb{E}_{\hat{\nu}_{n}*\Phi}[e^{f}]+1,$ (60a)
$\mathsf{D}_{\mathsf{KL}}\left(\mu*\Phi\middle\|\nu*\Phi\right)=\sup_{f\in\mathcal{F}_{n}}\mathbb{E}_{\mu*\Phi}[f]-\mathbb{E}_{\nu*\Phi}[e^{f}]+1.$ (60b)
Moreover, since $\left|f_{\mu*\Phi,\nu*\Phi}^{\star}(x)\right|\leq F_{n}(x)$ for all $x\in\mathbb{R}^{d}$ and $\left\|F_{n}*\phi\right\|_{p,\mu}<\infty$ for some $p\geq 2$ implies that $\left\|F_{n}\right\|_{1,\mu*\Phi}=\left\|F_{n}*\phi\right\|_{1,\mu}<\infty$, $\mathsf{D}_{\mathsf{KL}}\left(\mu*\Phi\middle\|\nu*\Phi\right)<\infty$. Then, from (44) and (60), we obtain
$\big{|}\mathsf{D}_{\mathsf{KL}}\left(\hat{\mu}_{n}*\Phi\middle\|\hat{\nu}_{n}*\Phi\right)-\mathsf{D}_{\mathsf{KL}}\left(\mu*\Phi\middle\|\nu*\Phi\right)\big{|}$
$\leq\sup_{f\in\mathcal{F}_{n}}\left|(\hat{\mu}_{n}*\phi-\mu*\phi)(f)-(\hat{\nu}_{n}*\phi-\nu*\phi)(e^{f})\right|$
$\leq\sup_{f\in\mathcal{F}_{n}}\left|(\hat{\mu}_{n}*\phi-\mu*\phi)(f)\right|+\sup_{f\in\mathcal{F}_{n}}\left|(\hat{\nu}_{n}*\phi-\nu*\phi)(e^{f})\right|$
$=\frac{Z_{n,\mathcal{F}_{n,\phi}}+\tilde{Z}_{n,\mathcal{E}_{n,\phi}}}{n},$